# CMB component-separated power spectrum estimation by Spectral Internal Linear Combination (SpILC)

### v1

@@ I Introduction
In order to study the wealth of signals in intensity and polarization maps across frequency channels, one is motivated to isolate the signals from each other. This includes separating extragalactic signals (e.g. CMB, tSZ, kSZ, cosmic infrared background, radio point sources) from each other and from galactic (e.g. dust, synchrotron) emissions. Various component separation methods are devised for this purpose, whereby the different frequency dependencies (spectral energy distributions, or SEDs) and spatial properties of components are exploited to produce a map of the desired component. These methods can be broadly separated into parametric and “blind” methods. Parameteric methods include Commander [32, 33, 34], which samples the joint posterior distribution of spectral parameters and component maps using a Bayesian approach. “Blind” methods are model agnostic except for the component of interest. Examples are SEVEM [35, 36, 37], which subtracts off a linear combination of internally-constructed foreground templates from a map to form a minimum-variance CMB map, and Internal Linear Combination (ILC) methods [38, 39, 40, 41], which also assume that the signal SED is known, and again seek a minimum-variance component map by a linear combination of maps across frequency channels. The “semi-blind” constrained ILC (cILC) [42] method additionally assumes the SEDs of a few modeled foreground components.
@@
where $x^{i}_{p}$ is the observed map at frequency channel $i$, $s_{p}$ is the component map one wishes to recover, $y_{p}$ is a modeled component map one wishes to explicitly remove, and $n^{i}_{p}$ are noise maps at channel $i$. The label $p$ denotes pixels, spherical harmonics or needlets for pixel-space, harmonic or needlet ILC, respectively. The SEDs $a^{i}$ and $b^{i}$ are assumed to be known exactly.111 In the case of the CIB where the SED is not exactly known, the moment ILC method [49] is found to be an effective mitigation [50] (discussion in Sec. VI).
@@ IV Data-split SpILC and cSpILC
This estimator will have lower variance than the estimator $\hat{K}^{\text{DS2}}_{\ell,11}=W_{\ell}^{ij}\hat{C}^{ij}_{\ell,12}$, as the construction of the latter discarded the information in the same-split, different-channel spectra $(1-\delta_{ij})\hat{C}^{ij}_{\ell,11}$ and $(1-\delta_{ij})\hat{C}^{ij}_{\ell,22}$. We note the estimator $\hat{K}^{\text{DS2}}_{\ell}$ instead of $\hat{K}^{\text{DS}}_{\ell}$ is required to remove bias from correlated atmospheric noise across channels, which is currently ignored in this study (we set the ensemble noise cross-spectra to be zero, $N^{ij}=0$ for $i\neq j$). The weights $W_{\ell}^{ij}$ are determined as usual by minimizing subject to constraints the variance of $\hat{K}_{\ell}$, which is derived in App. C.3.
@@ VII Acknowledgment
JK thanks Anthony Challinor, Steven Gratton, Selim Hotinli, Matthew Johnson, Terry Lam, Mathieu Remazeilles, and Kendrick Smith for discussions. We thank Mathieu Remazeilles for a careful reading of the manuscript, and Anthony Challinor for a careful reading of an earlier draft. JK acknowledges support from The Joseph Needham Foundation for Science & Civilisation (Hong Kong) through the Joseph Needham PhD fellowship. JK is grateful for the hospitality of Perimeter Institute for Theoretical Physics, where a part of this work was done. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ C.1 Data-split ILC and cILC
Suppose $x^{i}_{p}$ is a map constructed from data collected from time $t_{0}$ to $t_{0}+\Delta t$. Two noisier maps $x^{i,1}_{p}$ and $x^{i,2}_{p}$, referred to as data splits, can be constructed from the time segments $t_{0},t_{0}+\Delta t/2]$ and $t_{0}+\Delta t/2,t_{0}+\Delta t]$ respectively. The two time segments are labeled 1 and 2 respectively. The motivation of this construction is that noise in different data splits are assumed to be uncorrelated.
To eliminate the noise bias present in the auto-spectra of ILC and cILC maps, Ref. [53] proposes to produce ILC (cILC) maps $\hat{s}^{1}_{p}\equiv w^{i}_{\ell}x^{i,1}_{p}$ and $\hat{s}^{2}_{p}\equiv w^{i}_{\ell}x^{i,2}_{p}$ where weights are determined by minimizing the quantity
$\hat{\sigma}^{2}_{12}\equiv\frac{1}{N_{p}}\sum_{p}\hat{s}^{1}_{p}\hat{s}^{2}_{p}\;,$ (76)
subject to the corresponding constraints. The auto-spectrum of the signal is then estimated by $\hat{\sigma}^{2}_{12}$. Reference [53] claims that the weights $w_{\ell}^{i}$ has the same functional form as their ILC or cILC counterparts, with $\hat{C}_{\ell}^{ij}$ replaced with $\hat{C}^{ij}_{\ell,12}$, defined as
$\hat{C}^{ij}_{\ell,12}\equiv\frac{1}{N_{p}}\sum_{p}x^{i,1}_{p}x^{j,2}_{p}\;.$ (77)
The elimination of the noise bias can be simply demonstrated:
$\langle\hat{\sigma}^{2}_{12}\rangle$ $=\frac{1}{N_{p}}\sum_{p}\langle(s_{p}+w_{\ell}^{i}n^{i}_{p,1})(s_{p}+w_{\ell}^{j}n^{j}_{p,2})\rangle$
$=C^{ss}_{\ell}+w_{\ell}^{i}w_{\ell}^{j}\frac{1}{N_{p}}\sum_{p}\langle n^{i}_{1}n^{j}_{2}\rangle=C^{ss}_{\ell}\;.$ (78)
C.2 Optimality of the data-split ILC and cILC weights
While replacing $\hat{C}_{\ell}^{ij}$ with $\hat{C}^{ij}_{\ell,12}$ in the ILC and cILC weights seem to be the obvious modification to incorporate data splits, on closer inspection such weights can be seen only to maximize the (negative) ILC biases:
$\frac{\partial}{\partial w_{\ell}^{i}}\hat{\sigma}^{2}_{12}$ $=\frac{\partial}{\partial w_{\ell}^{i}}\frac{1}{N_{p}}\sum_{p}(s_{p}+w_{\ell}^{j}n^{j}_{p,1})(s_{p}+w_{\ell}^{k}n^{k}_{p,2})$
$=\frac{\partial}{\partial w_{\ell}^{i}}\frac{1}{N_{p}}\sum_{p}\left[w_{\ell}^{j}s_{p}(n^{j}_{p,1}+n^{j}_{p,2})+w_{\ell}^{j}w_{\ell}^{k}n^{j}_{p,1}n^{k}_{p,2}\right]\;.$ (79)
In other words, if the ILC biases vanishes or if we determine the weights with true covariances, any set of weights that satisfies the constraints would recover the signal map $s_{p}$ identically.
However, using an arbitrary set of weights would not minimize the variance of $\hat{\sigma}^{2}_{12}$ :
@@ C.2 Optimality of the data-split ILC and cILC weights
$=\frac{2}{N_{p}}\left((C^{ss}_{\ell})^{2}+\frac{1}{2}w_{\ell}^{j}w_{\ell}^{k}w_{\ell}^{m}w_{\ell}^{n}(2N_{\ell}^{jm})(2N_{\ell}^{kn})\right)$
$=\frac{2}{N_{p}}(C^{ss}_{\ell})^{2}+\frac{1}{N_{p}}\left[2w_{\ell}^{j}w_{\ell}^{k}N_{\ell}^{jk}\right]^{2}\;,$ (80)
@@
The weights above which only maximizes the (negative) ILC biases does not minimize the quantity $w_{\ell}^{j}w_{\ell}^{k}N_{\ell}^{jk}$, thus do not minimize the variance of the signal auto-spectrum estimator $\hat{\sigma}^{2}_{12}$.
The weights that actually minimize $w_{\ell}^{j}w_{\ell}^{k}(\sigma^{2}_{n})^{jk}$ are actually the normal ILC and cILC weights, which minimizes subject to constraints the (ensemble average of) the total error variance:
$\frac{\partial}{\partial w_{\ell}^{i}}\chi^{2}\equiv\frac{\partial}{\partial w_{\ell}^{i}}\langle(\hat{s}_{p}-s_{p})^{2}\rangle=0\;,$
$\implies\frac{\partial}{\partial w_{\ell}^{i}}w_{\ell}^{j}w_{\ell}^{k}\langle n_{p}^{j}n^{k}_{p}\rangle=\frac{\partial}{\partial w_{\ell}^{i}}w_{\ell}^{j}w_{\ell}^{k}N_{\ell}^{jk}=0\;.$ (82)
@@ C.3 Variance of Data-Split Estimators
Consider the data-split estimator in Eq. (34),

### v2

@@ I Introduction
In order to study the wealth of signals in intensity and polarization maps across frequency channels, one is motivated to isolate the signals from each other. This includes separating extragalactic signals (e.g. CMB, tSZ, kSZ, cosmic infrared background, radio point sources) from each other and from galactic (e.g. dust, synchrotron) emissions. Various component separation methods are devised for this purpose, whereby the different frequency dependencies (spectral energy distributions, or SEDs) and spatial properties of components are exploited to produce a map of the desired component. These methods can be broadly separated into parametric and “blind” methods. Parameteric methods include Commander [20, 21, 40], which samples the joint posterior distribution of spectral parameters and component maps using a Bayesian approach. “Blind” methods are model agnostic except for the component of interest. Examples are SEVEM [36, 30, 22], which subtracts off a linear combination of internally-constructed foreground templates from a map to form a minimum-variance CMB map, and Internal Linear Combination (ILC) methods [19, 60, 15, 37], which also assume that the signal SED is known, and again seek a minimum-variance component map by a linear combination of maps across frequency channels. The “semi-blind” constrained ILC (cILC) [50] method additionally assumes the SEDs of a few modeled foreground components. At the power spectrum level, the Analytical Blind Separation (ABS) method [64, 63] performs eigenmode filtering on the observed cross spectra, building upon a unique solution for the signal power in the zero-noise case with more channels than foreground components.
@@
where $x^{i}_{p}$ is the observed map at frequency channel $i$, $s_{p}$ is the component map one wishes to recover, $y_{p}$ is a modeled component map one wishes to explicitly remove, and $n^{i}_{p}$ are noise maps at channel $i$. The label $p$ denotes pixels, spherical harmonics or needlets for pixel-space, harmonic or needlet ILC, respectively. The SEDs $a^{i}$ and $b^{i}$ are assumed to be known exactly. [^1] In the case of the CIB where the SED is not exactly known, the moment ILC method [10] is found to be an effective mitigation [38] (discussion in Sec. VI).
@@ IV Data-split SpILC and cSpILC
This estimator will have lower variance than the estimator $\hat{K}^{\text{DS2}}_{\ell,11}=W_{\ell}^{ij}\hat{C}^{ij}_{\ell,12}$, as the construction of the latter discarded the information in the same-split, different-channel spectra $(1-\delta_{ij})\hat{C}^{ij}_{\ell,11}$ and $(1-\delta_{ij})\hat{C}^{ij}_{\ell,22}$. We note the estimator $\hat{K}^{\text{DS0}}_{\ell}$ instead of $\hat{K}^{\text{DS}}_{\ell}$ is required to remove bias from correlated atmospheric noise across channels, which is currently ignored in the variance comparison (in simulations we set the ensemble noise cross-spectra to be zero, $N^{ij}=0$ for $i\neq j$). The weights $W_{\ell}^{ij}$ are determined as usual by minimizing subject to constraints on the variance of $\hat{K}_{\ell}$, which is derived in App. C.3 for both $\hat{K}^{\text{DS0}}_{\ell}$ and $\hat{K}^{\text{DS}}_{\ell}$.
@@ VII Acknowledgment
JK thanks Anthony Challinor, Steven Gratton, Colin Hill, Selim Hotinli, Matthew Johnson, Terry Lam, Mathieu Remazeilles, and Kendrick Smith for discussions. We thank Mathieu Remazeilles for a careful reading of the manuscript, and Anthony Challinor for a careful reading of an earlier draft. JK acknowledges support from The Joseph Needham Foundation for Science & Civilisation (Hong Kong) through the Joseph Needham PhD fellowship. JK is grateful for the hospitality of Perimeter Institute for Theoretical Physics, where a part of this work was done. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ C.1 Data-split ILC and cILC
Suppose $x^{i}_{p}$ is a map constructed from data collected from time $t_{0}$ to $t_{0}+\Delta t$. Two noisier maps $x^{i,1}_{p}$ and $x^{i,2}_{p}$, referred to as data splits, can be constructed from the time segments $t_{0},t_{0}+\Delta t/2]$ and $t_{0}+\Delta t/2,t_{0}+\Delta t]$ respectively. The two time segments are labeled 1 and 2 respectively. The noise in different data splits are assumed to be uncorrelated.
An ILC method built on these splits was proposed in Ref. [26], where the original motivation is to minimize the foreground variance in the auto-spectra of ILC and cILC maps by preventing instrumental noise from contributing to the weights. Here we review this approach, show that it is not the minimum-variance power spectrum estimator, and derive the minimum-variance data-split estimator using the SpILC method proposed in this work.
Ref. [26] proposes to produce ILC (cILC) maps $\hat{s}^{1}_{p}\equiv w^{i}_{\ell}x^{i,1}_{p}$ and $\hat{s}^{2}_{p}\equiv w^{i}_{\ell}x^{i,2}_{p}$ where weights are determined by minimizing the sample covariance between two different splits
$\hat{\sigma}^{2}_{12}\equiv\frac{1}{N_{p}}\sum_{p}\hat{s}^{1}_{p}\hat{s}^{*,2}_{p}\;,$ (76)
subject to ILC (or cILC) constraints. The auto-spectrum of the signal is then estimated by $\hat{\sigma}^{2}_{12}$. The elimination of the noise bias can be simply demonstrated:
$\langle\hat{\sigma}^{2}_{12}\rangle$ $=\frac{1}{N_{p}}\sum_{p}\langle(s_{p}+w^{i}_{\ell}k^{i}_{p}+w_{\ell}^{i}n^{i}_{p,1})(s^{*}_{p}+w^{j}_{\ell}k^{*,j}_{p}+w_{\ell}^{j}n^{*,j}_{p,2})\rangle$
$=C^{ss}_{\ell}+w^{i}_{\ell}w^{j}_{\ell}\frac{1}{N_{p}}\sum_{p}\langle k^{i}_{p}k^{*,j}_{p}\rangle+w_{\ell}^{i}w_{\ell}^{j}\frac{1}{N_{p}}\sum_{p}\langle n^{i}_{p,1}n^{*,j}_{p,2}\rangle$
$=C^{ss}_{\ell}+w^{i}_{\ell}w^{j}_{\ell}F^{ij}_{\ell}\;,$ (77)
where $k^{i}_{p}$ is the foreground map at the $i$-th frequency, and $F^{ij}_{\ell}\equiv\langle k^{i}_{p}k^{j}_{p}\rangle$ is the foreground cross spectra. Here we assumed that foregrounds are independent from the CMB and instrumental noise. If cILC constraints are imposed, then the foreground term $w^{i}_{\ell}w^{j}_{\ell}F^{ij}_{\ell}$ can be further reduced or eliminated at a noise cost.
C.2 Optimal weights for the data-split ILC and cILC power spectrum estimators
Constructing data splits that use weights that minimizes $\hat{\sigma}^{2}_{12}$ minimizes the foreground sample covariance $\sum_{p}w^{j}_{\ell}w^{k}_{\ell}k_{p}^{j}k_{p}^{k*}/N_{p}$ together with sample covariance due to chance correlations between signal, foreground and noise in the two splits. If cILC constraints are imposed, even the foreground sample covariance vanishes, and the weights only minimizes these chance correlations i.e. ILC biases.
Instead, if one wishes to minimize the variance of the estimated auto power spectrum $\hat{\sigma}^{2}_{12}$, the weights discussed above is not optimal. Writing out the variance:
@@ C.2 Optimal weights for the data-split ILC and cILC power spectrum estimators
$=\frac{1}{N_{p}}\left(C^{ss}_{\ell}+w^{j}_{\ell}w^{k}_{\ell}F^{jk}_{\ell}\right)^{2}$
$\phantom{{}={}}+\frac{1}{N_{p}}\left(C^{ss}_{\ell}+w^{j}_{\ell}w^{k}_{\ell}F^{jk}_{\ell}+2w_{\ell}^{j}w_{\ell}^{k}N_{\ell}^{jk}\right)^{2}\;,$ (78)
@@
The weights above which neglects the noise power $w^{i}_{\ell}w^{j}_{\ell}N^{ij}_{\ell}$ does not minimize the variance of the signal auto-spectrum estimator, $\operatorname{Var}(\hat{\sigma}^{2}_{12})$.
On the other hand, note that the normal ILC and cILC weights (using the ensemble covariance) minimize (subject to constraints)
$\operatorname{Var}(\hat{s}^{\text{ILC}})$ $=C^{ss}_{\ell}+w^{j}_{\ell}w^{k}_{\ell}(F^{jk}_{\ell}+N^{jk}_{\ell})\;,$ (80)
$\operatorname{Var}(\hat{s}^{\text{cILC}})$ $=C^{ss}_{\ell}+w^{j}_{\ell}w^{k}_{\ell}N^{jk}_{\ell}\;.$ (81)
One therefore sees that the cILC weights are the optimal weights for minimizing the variance $\operatorname{Var}(\hat{\sigma}^{2}_{12})$ of the cILC data-split power spectrum. However, neither the ILC nor the ILC data-split weights above minimizes $\operatorname{Var}(\hat{\sigma}^{2}_{12})$. To build a data-split power spectrum estimator with minimum variance one has to use the SpILC formalism, discussed in the next section.
@@ C.3 Variance of Data-Split Estimators
Depending on the observational setup, the noise may be independent across different frequency channels or may have correlations across frequencies (which is the case for e.g. atmospheric noise). Here we propose two data-split power spectrum estimators: the first (DS0) uses only cross spectra between splits and is applicable when the noise is correlated across frequency channels, and the second (DS, considered in the main text) discards only the split auto spectra (same split and same frequency) and is applicable for the independent noise case.
The first estimator generalizes the estimator in the previous sub-section:
$\hat{K}_{\ell}^{\text{(DS0)}}=\sum_{ij}W^{ij}_{\ell}\hat{C}_{\ell,12}^{ij}\;,$ (82)
the variance is
$\operatorname{Var}(\hat{K}_{\ell}^{\text{(DS0)}})$ $=\frac{2}{2\ell+1}W^{jk}_{\ell}W^{mn}_{\ell}C^{jm}_{\ell,1(1|}C^{kn}_{\ell,2|2)}\;.$ (83)
$\hat{C}_{\ell}^{(m|i}\hat{C}_{\ell}^{|n)j}\mapsto$ $\hat{C}^{{\color[rgb]{1,0,0}(m_{\alpha}|}{\color[rgb]{0,0,1}(i_{\beta}|}}_{\ell,(1|1}\hat{C}^{{\color[rgb]{1,0,0}|n_{\alpha})}{\color[rgb]{0,0,1}|j_{\beta})}}_{\ell,|2)2}\;$. (84)
These symmetrization brackets are necessary to ensure that the object is symmetric in $j\leftrightarrow k$ and $m\leftrightarrow n$ during the optimization. Once this is substituted in the $D^{\mu\nu}_{\ell}$ matrix and Eq. (56) is solved, one gets the weights for $\hat{K}^{(\text{DS0})}_{\ell}$.
Now consider the second data-split estimator, which is considered in the main text in Eq. (34), where only same-split, same-frequency information is discarded: